Is (-7, -9) a solution of
2x -4y =-6

Answer by stanbon(75887)

Is (-7, -9) a solution of
2x -4y =-6
---
Substitute:
2*-7 -4*-9 = -6
-14+36 = -6
False.
Ans: No
==============
Cheers,
Stan H.